ACIS Thermal/Vacuum Test Page

Latest update: Sun Apr 13 16:22:12 1997 EST

ACIS Home Page Status Plan Changes Events Notes Anomalies Phone Numbers Shifts
Live Engineering Data Phase II Plan Procedures
Thermal/Vacuum Long-Form Short-Form

Major Hardware Changes

  • Noisy PSMC telemetry encountered in first TV was due to LRCTU which still had the OFFICIAL RCTU serial digital interface; we fixed it.
  • The science data noise came from two sources -- DPA/DEA crosstalk and internal power line noise; fixes have been applied for both problems.
  • Fixes were applied to the Focal Plane thermal system; we should now be able to achieve -120C reliably.
  • The optical light leak between the two Optical Blocking Filters has been taped.
  • A harness change in the DEA enabled the OBF-test LED.

    Current Status:

  • In the shipping box.

    Short Term Plan:


    Test Event History

    See the
    Phase II Plan for references to the event descriptions.
    Day Planned Event Plan Ref. Actual Events Results
    4/6, Sun Short Form/Initial Pump 1,2
    SF OK; start pumping 2200
    4/7, Mon Baseline Performance 3 Took HST data on all CCDs They are all there; BI!=FI

    Initial Bake 4 Start at 10:30am
    4/8, Tue LN2 into shrouds 4b,5,6 Cool to ambient
    Vent
    TQCM 2Hz/Hr since Midnight; Chamber pressure 4x10(-6)
    Low Conductance Valve opened at 10am.

    Cool shrouds 4a,7,8,9,10,11,12 High Conductance Valve opened at 12noon
    LN2 into shroud at 1:30pm
    Major heertburn; see Anomaly section

    Door Tests 8,9,10,11,12 Door Tests ran fine Drop DH bakeout mode at 1700
    4/9, Wed. Science Cold 13,14
    FP reached -120 at 7am.
    Short Form OK.
    Internal LED and Fe55 source OK.
    4/10, Thu. Science Hot 15,16,17,18 Long Form Science good; Long Form OK
    4/11, Fri. 1238 Cert. 21,22,23
    Cert. started at 11:15am
    4/12, Sat. Remove from Chamber 24,25,26,27
    Large Valve was open


    Test Notes

    • The flight CCDs, run in diagnostic mode at room temperature, seem to be running bias-subtracted sigmas of 2.0-2.5 ADU rms. When cold, and analyzed by real scientists, the results are 2.2-2.7 ADU resolution. When the electronics are warm, the results are 1.8-2.2 ADU.
    • Backside Illuminated CCDs aren't like Frontside Illuminated CCDs, but this was the Version 1.2 load of the Flight Software with the analog board parameters set for all Frontside chips. When run with a hand load of the 1.3 parameters, these chips show the same initial 100ADU change in the first 20 colums, but the vertical variations disappear.
      See /nfs/glenlivet/d4/capture/ffp[is][05].*.fits for ambient diag. data.
    • Running in cold soak, we're burning about 1.8w keeping the focal plane at -120C, and about 10w keeping the Detector Housing at -60C. (That's about 45% of max power for the focal plane heater; about 25% of max power for the DH heater -- but it's also used for bake out.)
      In hot soak there is none to negative margin.

      Anomalies, Major and Minor:

      • On the Engineering Unit DH the low conductance valve took 1 hour to reduce internal pressure from 14 torr to 1.5 torr; the Flight unit will take a fair amount longer.
        On the Flight Unit DH the low conductance valve took 1 hour to reduce internal pressure from 18 torr to 3.25 torr; 2 hours to achieve 2.0 torr. We should have waited >3 hours before opening Large Valve if we use the 1 torr max. criteria.
      • Through major operator error (RFG) managed to get the Vent Valve Actuators fighting each other, resulting in an ambiguous final state from which neither open nor close commands will complete. This will require some serious hardware inspection.
      • Troubleshooting of the large vent valve actuator produced no new information and the chamber was repressurized at 5 torr/minute with the valve open (as indicated by the DH pressure transducers failing to show any hint of a differential pressure).
      • Upon inspection of the actuator it was seen that one shear disk had ruptured. The actuator was replaced with a flight spare and checked out using the Vacuum GSE.

        Phone Numbers

        (from MIT, Lincoln Lab is 181 + 4 digits)

        (from Lincoln Lab, MIT is 806 + 5 digits)

        Name Function Work Home Notes
        Ellen Sen Ultimate TC 253-8342 617-665-3568
        Demitrios Athens Software GSE 258-5285 617-566-7668
        Mark Bautz Proj. Science 253-0023 617-863-2024
        Mike Doucette Electrical GSE 253-3571 617-438-7690
        Bob Goeke Proj. Eng. 253-1910 617-729-8484
        Brian Klatt Quality 253-7555 508-785-1511
        Bill Mayer Proj. Mgr. 253-7552 508-692-3318
        Matt Smith Contamination 252-1736 508-435-9970
        Ed Boughan TC 253-6146 508-470-1043
        Mary Briggs TC 258-7896 508-372-0358
        Sean Callaghan TC
        617-969-6178
        Peter Ford TC 253-7277 617-661-0868
        Gordon Gong TC 253-8135 617-969-3768
        Dorothy Gordon TC 253-0797 617-232-7327
        Hans Govaert TC 253-1977 508-535-1467
        John Hughes TC 253-8009 508-497-8344
        Fred Kasparian TC 253-1065 617-327-6330
        Jim O'Connor TC 253-6139 617-335-2915
        Joanne Vining TC 258-8090 617-396-9737
        Engineering GSE TC sits here 981-4604

        Science GSE
        981-2010

        Cleanroom/TV
        981-5138

        Ron Efromson LL Coordinator 981-4513

        Mike Wirzbicki LL Contamination 981-4326 981-5348 (beep)617-430-1886
        Jon Howell LL Vacuum 981-4990
        (beep)617-430-5684
        Angelo Calao LL Manager 981-7717


        Engineering GSE Shifts

        Day 12mid-8am 8am-4pm 4pm-12mid
        4/7, Monday
        Sen Mayer
        4/8, Tuesday Callaghan Boughan Gong
        4/9, Wednesday Callaghan Vining Cumings
        4/10, Thursday Callaghan Briggs Kasparian
        4/11, Friday Callaghan Vining Klatt
        4/12, Saturday Callaghan Kasparian Mayer
        4/13, Sunday Callaghan Gordon Klatt


        Previous as-run T/V test record is still available.

        Page maintainer: Bob Goeke.

        There have been 27 accesses to this page since 9/17/2022.